# Peer-induced Fairness: A Causal Approach for Algorithmic Fairness Auditing

### v3

@@ 4.1 Discrimination from peer comparisons
The phenomenon of discrimination, a ubiquitous aspect of daily life, is extensively explored within cognitive science literature. Research indicates that perceptions of discrimination are shaped not only by personal experiences but also through comparisons with peers who, despite possessing similar capabilities, skills, or knowledge, experience differential treatment, leading to missed opportunities. These perceptions are cultivated both through individual encounters and the lens of peer experiences. When an individual’s treatment aligns with that of their peer group, perceptions of being biased tend to diminish. Studies have shown that social and financial ties are more likely to form among individuals who share similarities in revenue levels, consumption behaviours, educational background, class, gender, race, or creditworthiness, illustrating a preference for homogeneity (Li et al.,, 2020; Haenlein,, 2011; Goel and Goldstein,, 2014; Wei et al.,, 2016).
@@ 4.2 Fairness through peer observations
Consider an individual $A$ from a protected group with a protected status $S=s_{-}$ and other unprotected features $\bm{X}=\bm{x}$, denoted as $A=(s_{-},\bm{x})$. Assuming the protected and unprotected groups are comparable, if there exists a group of peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ from the unprotected group $S=s_{+}$, represented as $\{(s_{+},\bm{x}_{1}),(s_{+},\bm{x}_{2}),\cdots\}$, forming an $A$-oriented network. We use the expectation of the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})$ across these peers $\mathcal{C}$ to approximate the counterfactual $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$, mathematically expressed as
@@ 4.3 Peer definition and identification
Before initiating peer comparisons, we need to formulate the definition of peers.
@@
Let us consider an individual $A$ belonging to a protected group, characterised by a protected feature $S=s_{-}$ and a set of unprotected features $\bm{X}=\bm{x}_{0}$, represented as $A=(s_{-},\bm{x}_{0})$. Assuming there exists a set of individuals =$\mathcal{B}=\{B_{1},B_{2},\cdots\}$ from the unprotected group, where $B_{i}=(s_{+},\bm{x}_{i})$ for $i=1,2,\ldots$. An individual $C\in\mathcal{B}$ is defined as $\delta$-peer of $A$ if the difference in joint distributions between $C$’s actual SWIG, $\mathcal{G}(s_{+},\bm{x}_{j})$, and $A$’s counterfactual SWIG, $\mathcal{\tilde{G}}(s_{+},\bm{x}_{0})$, is less than a threshold $\delta$,
@@
The concept of a peer in the graphical causal model is defined through the interrelations among three random variables: $S$, $\bm{X}$, and $Y$. For rigorous and unbiased comparisons, it is essential that a peer exhibits a joint distribution similar to the counterfactual scenario.
Despite the appealing theoretical foundation of the $\delta$-peer concept, its practical implementation encounters significant challenges. A primary obstacle is the difficulty in calculating $\mathbb{P}(Y(s_{+},\bm{x}))$ from Eq. (6) for the counterfactual case $(s_{+},\bm{x})$, which is crucial for assessing peer similarity in such contexts. This complication stems from the representation of $\bm{x}$ as the unprotected features for the protected group, where direct calculation of this probability is often unfeasible due to the absence of observational data. To address this and develop a more feasible approach for peer selection, we re-examine Eq. (3) and Eq. (6).
Since it is not feasible to directly derive $\mathcal{\tilde{G}}(s_{+},\bm{x})$ from observational data, we have no choice but use the information from $\mathcal{G}(s_{+},\bm{x})$ as a proxy for approximation, which has been discussed in Section 4.2. Upon comparing Eq. (3) and Eq. (6), the difference lies in the terms $\bm{X}$ and $Y$. Referring to Panel (a) of Fig. 1 and considering $\bm{x}_{0}$ as the observable unprotected features of an individual from the protected group $S=s_{-}$, we can compute $\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ using Bayes’ formula:
@@
However, because $\bm{x}_{0}$ are the observable unprotected features for an individual from the protected group $S=s_{-}$, estimating $\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ directly is not feasible. Given that $S$ represents a binary set, we can infer:
$\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ $=1-\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}_{0}),$ (11)
$\mathbb{P}(S=s_{+})$ $=1-\mathbb{P}(S=s_{-}).$ (12)
We propose a unified notation for both $\mathbb{P}(\bm{X}(s_{-}))$ and $\mathbb{P}(\bm{X}(s_{+}))$ to streamline calculations and ensure consistency across analyses:

### v4

@@ 3.1 Discrimination from peer comparisons
The phenomenon of discrimination, a ubiquitous aspect of daily life, is extensively explored within cognitive science literature. Research indicates that perceptions of discrimination are shaped not only by personal experiences but also through comparisons with peers who, despite possessing similar capabilities, skills, or knowledge, experience differential treatment, leading to missed opportunities. These perceptions are cultivated both through individual encounters and the lens of peer experiences (Li and Jain,, 2016). When an individual’s treatment aligns with that of their peer group, perceptions of being biased tend to diminish. Studies have shown that social and financial ties are more likely to form among individuals who share similarities in revenue levels, consumption behaviours, educational background, class, gender, race, or creditworthiness, illustrating a preference for homogeneity (Li et al.,, 2020; Haenlein,, 2011; Goel and Goldstein,, 2014; Wei et al.,, 2016).
The insights from cognitive science highlight the importance of understanding how comparative experiences shape perceptions of discrimination among individuals and organisations. These groups may not only perceive but also actually experience biased outcomes when compared to their peers. Such perceived or real biases could erode trust in automated systems and, more broadly, undermine confidence in the regulators intended to ensure fairness. To mitigate these risks, regulators need to ensure that auditing frameworks are sensitive to these subtler forms of discrimination, which can arise from differences in treatment relative to peers.
@@ 3.2 Fairness through peer observations
Consider an individual $A$ from a protected group with a protected status $S=s_{-}$ and other unprotected features $\bm{X}=\bm{x}$, denoted as $A=(s_{-},\bm{x})$. Assuming the protected and unprotected groups are comparable, if there exists a group of peers $\mathcal{C}=\{C_{1},C_{2},\cdots\}$ from the unprotected group $S=s_{+}$, represented as $\{(s_{+},\bm{x}_{1}),(s_{+},\bm{x}_{2}),\cdots\}$, forming an $A$-oriented network. We use the expectation of the probability $\mathbb{P}(\hat{Y}_{s_{+}}|s_{+},\bm{x}_{j})$ across these peers $\mathcal{C}$ to approximate the counterfactual $\mathbb{P}(\hat{Y}_{s_{+}}|s_{-},\bm{x})$, mathematically expressed as:
@@ 3.3 Peer definition and identification
While the counterfactual predictive probability is provided in Eq. (7), a key question remains: “What is a peer” in mathematical terms? A precise definition and identification of peers are crucial before initiating peer comparisons, as they ensure accurate and reliable assessments. This clarity is essential for regulators to effectively audit potential biases in algorithmic decision-making.
@@
Let an individual $A$ belong to a protected group, characterised by a protected feature $S=s_{-}$ and a set of unprotected features $\bm{X}=\bm{x}_{0}$, represented as $A=(s_{-},\bm{x}_{0})$. Assuming there exists a set of individuals =$\mathcal{B}=\{B_{1},B_{2},\cdots\}$ from the unprotected group, where $B_{i}=(s_{+},\bm{x}_{i})$ for $i=1,2,\ldots$. An individual $C\in\mathcal{B}$ is defined as $\delta$-peer of $A$ if the difference in joint distributions between $C$’s actual SWIG, $\mathcal{G}(s_{+},\bm{x}_{j})$, and $A$’s counterfactual SWIG, $\mathcal{\tilde{G}}(s_{+},\bm{x}_{0})$, is less than a threshold $\delta$,
@@
In a graphical causal model, the concept of a peer is defined through the interrelations among three random variables: $S$, $\bm{X}$, and $Y$. To enable rigorous and unbiased comparisons, it is crucial that a peer exhibits a joint distribution similar to the counterfactual scenario. However, since the counterfactual scenario is inherently unobservable, it is rarely possible to observe the exact same $\bm{X}$ and $Y$ in another group defined by different protected attributes.
Although we have the mathematical representation of a $\delta$-peer in Definition 3, its practical implementation faces significant challenges. A major obstacle is the difficulty in calculating $\mathbb{P}(Y(s_{+},\bm{x}))$ from Eq. (6) for the counterfactual scenario $(s_{+},\bm{x})$, which is essential for evaluating peer similarity. This difficulty arises because $\bm{x}$ represents the unprotected features for the protected group, where the direct calculation of this probability is often infeasible due to the lack of observational data.
To address this and develop a more feasible approach for peer selection, we re-examine Eq. (3) and Eq. (6). Since it is not feasible to directly derive $\mathcal{\tilde{G}}(s_{+},\bm{x})$ from observational data, we have no choice but use the information from $\mathcal{G}(s_{+},\bm{x})$ as a proxy for approximation, which has been discussed in Section 3.2. Upon comparing Eq. (3) and Eq. (6), the difference lies in the terms $\bm{X}$ and $Y$. Referring to Figure 1 (a) and considering $\bm{x}_{0}$ as the observable unprotected features of an individual from the protected group $S=s_{-}$, we can compute $\mathbb{P}(\bm{X}(s_{-})=\bm{x}_{0})$ using Bayes’ formula:
@@
However, because $\bm{x}_{0}$ are the observable unprotected features for an individual from the protected group $S=s_{-}$, estimating $\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})$ directly is not feasible. Given that $S$ represents a binary set, we can infer $\mathbb{P}(S=s_{+}|\bm{X}=\bm{x}_{0})=1-\mathbb{P}(S=s_{-}|\bm{X}=\bm{x}_{0})$. We can rewrite $\mathbb{P}(\bm{X}(s_{-}))$ and $\mathbb{P}(\bm{X}(s_{+}))$ in a unified representation,